OpenStudy (anonymous):

integrate the following:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x^3 }{ \sqrt{x^8 + 1} } dx\]

OpenStudy (yrelhan4):

Its x^3 in the numerator right?

OpenStudy (jhannybean):

\[\bf\large \int\limits \frac{ x^3 }{ \sqrt{x^8+1} }\]

OpenStudy (anonymous):

let u = x^4. if you're familiar with substitution you should have this

OpenStudy (yrelhan4):

you forgot the dx... what you need to do is.. substitute x^4=t ==> x^8=t^2 and 4x^3dx=dt..

OpenStudy (jhannybean):

ik i forgot the dx-_- accident :(

OpenStudy (anonymous):

hahahah, thanks guys, my friend wanted to confirm this answer, so i just posted it here to check :)

OpenStudy (yrelhan4):

:P, Jhanny

OpenStudy (yrelhan4):

Cheers, Tanvidais. :)

OpenStudy (anonymous):

ok wait hold on.

OpenStudy (jhannybean):

just a question... you have \[\bf\large \int\limits \frac{ x^3 }{ \sqrt{(x^4)^{2}+1} }\] so cant you write the square root part as tan inverse?

OpenStudy (anonymous):

integrate this one too??: \[\int\limits_{}^{}\frac{ 1 }{ \sqrt{sinx} }dx\]

OpenStudy (jhannybean):

arg, no dx, they need equation editor -_-

OpenStudy (yrelhan4):

sinx= (2tanx/2)/1+tan^2 x/2, i guess? correct me if i ma wrong. i am not too sure if its 1 + .. or 1- ..

OpenStudy (yrelhan4):

I just googled.. it is 1+tan^2 x/2 in the denominator.. hmm. this could help if there wasnt the sqrt.. i am not too sure what next..

OpenStudy (jhannybean):

Where is THAT identity?? I've never seen \[\large \sin x = \frac{ 2tanx }{ 2 }*\frac{ 1 }{ 1+\tan ^{2}(\frac{ x }{ 2 }) }\]

OpenStudy (yrelhan4):

umm. sin2x = 2tanx/1+tan^2 x.. so just substitute x by x/2.. and its 2tan(x/2).. not 2tanx/2.. sorry if i mistyped earlier.

OpenStudy (jhannybean):

x_x i'm confused, what did you google.... I just want to see some one work out the identity :) I'd like to learn this :P

OpenStudy (yrelhan4):

There you go. :)

OpenStudy (yrelhan4):

and i dont think i can integrate the second one.

OpenStudy (jhannybean):

I atleast understand the sin 2x identity now:) haha

OpenStudy (yrelhan4):

http://www.wolframalpha.com/input/?i=integration+1%2Fsqrt+sinx Wolf says this.. if anyone can understand it. :P

OpenStudy (jhannybean):

what math is this??? lol. elliptical integral of the first kind?

OpenStudy (jhannybean):

hyperbolic identity?

OpenStudy (hba):

oh mimi is so intelligent :D

OpenStudy (yrelhan4):

eh we are done with the original question Mimi.. 1/sqrt(sinx) is the question. :P

OpenStudy (jhannybean):

\[\large \color{darkorchid}{sinhx=\frac{ e ^{x}-e ^{-x} }{ 2 }}\]

OpenStudy (yrelhan4):

umm. arent we done with x^3 dx/ sqrt ( x^8 + 1) ??

OpenStudy (jhannybean):

Yeah. She was just answering both questions

OpenStudy (yrelhan4):

Ohhh sorry. :P

OpenStudy (jhannybean):

This is for the second one? Looks more like the first one haha. Oh it is the first one...

OpenStudy (yrelhan4):

Lol Mimi.. we were done with the orginal question.. then he asked the second question.. 1/ sqrt sinx.. :P

OpenStudy (jhannybean):

We're all stuck on \[\large \int\limits \frac{ 1 }{ \color{green}{\sqrt{\sin x}} }dx\]

OpenStudy (jhannybean):

@.Sam.

sam (.sam.):

Let u=x^4 du=4x^3dx du/4=x^3dx \[=\frac{1}{4} \int\limits \frac{1}{\sqrt{1+u^2}} \, du\] Then use trig sub since a^2+x^2, we'll use \[u=\tan \theta \\ \\ du=\sec^2\theta d \theta\] \[=\frac{1}{4} \int\limits \frac{1}{\sqrt{1+u^2}} \, du\] \[=\frac{1}{4} \int\limits \frac{1}{\sqrt{1+tan^2 \theta}} sec^2\theta\ d\theta \] Using \(1+\tan^2 \theta=\sec^2\theta\), \[=\frac{1}{4} \int\limits \frac{1}{\sqrt{\sec^2\theta}} sec^2\theta\ d\theta \] \[=\frac{1}{4} \int\limits sec\theta\ d\theta \]

OpenStudy (yrelhan4):

SAM SAM SAM.. WE ARE DONE WITH THAT QUESTION. :P WE ARE STUCK ON A DIFFERENT QUESTION.. 1/sqrt (sinx) .. he asked that later in the thread..

sam (.sam.):

ew

OpenStudy (yrelhan4):

:P